# Maximum packings in graphs forbidding given rainbow cycles

### 22 Mar 2026 v1

@@ Abstract
Motivated by the Ruzsa-Szemerédi problem, Imolay, Karl, Nazy, and Váli studied a variant of Turán number $ex_{F}(n,G)$ (called the $F$-multicolor Turán number of $G$), defined as the maximum number of edge-disjoint copies of $F$ on $n$-vertex set such that there is no copies of $G$ whose edges come from distinct copies of $F$. They proved that if there is no homomorphism from $G$ to $F$, then $n^{2}/v(F)^{2}+o(n^{2})\leq ex_{F}(n,G)\leq ex(n,G)/e(F)+o(n^{2})$, and otherwise $ex_{F}(n,G)=o(n^{2})$. The quantity $ex_{F}(n,G)$ asymptotically equals the maximum size of an $F$-packing in an $n$-vertex $G$-free graph, and attains the upper bound $ex(n,G)/e(F)+o(n^{2})$ if and only if $\chi(G)>\chi(F)$.
@@ 1 Introduction
The $(6,3)$-problem can be equivalently stated as follows: determine the maximum number of edge-disjoint triangles in an $n$-vertex graph such that no triangle is formed by taking one edge from each of three distinct triangles. Imolay, Karl, Nazy and Váli [13] generalized this problem by considering the following: given two fixed graphs $F$ and $G$, determine the maximum number of edge-disjoint copies of $F$ in an $n$-vertex graph such that there is no copy of $G$ whose edges come from distinct copies of $F$. This maximum, denoted by $ex_{F}(n,G)$, is called the $F$-multicolor Turán number of $G$. Clearly, if $F$ and $G$ are $r$-uniform hypergraphs, then $ex_{F}(n,G)=O(n^{r})$.
@@
In what follows, we restrict our discussion to simple graphs. Imolay, Karl, Nazy and Váli [13] proved that $ex_{F}(n,G)=\Theta(n^{2})$ if and only if there is no homomorphism from $G$ to $F$, and for the nondegenerate case ($ex_{F}(n,G)=\Theta(n^{2})$),
@@
Li [15] extended the above results to uniform hypergraphs. We say $ex_{F}(n,G)$ attains its lower bound (resp. upper bound) in (1) if $ex_{F}(n,G)=n^{2}/v(F)^{2}+o(n^{2})$ (resp. $ex_{F}(n,G)=ex(n,G)/e(F)+o(n^{2})$). Note that $\chi(G)>\chi(F)$ implies that there is no homomorphism from $G$ to $F$. Imolay, Karl, Nazy and Váli [13] showed that $ex_{F}(n,G)$ attains its upper bound when $\chi(G)>\chi(F)$. In fact, using the stability of extremal graphs for $ex(n,G)$, one can show that this condition is also necessary: $ex_{F}(n,G)$ attains its upper bound if and only if $\chi(G)>\chi(F)$ [15].
Imolay, Karl, Nazy and Váli [13] posed the question of characterizing all graph pairs $(F,G)$ for which $ex_{F}(n,G)=n^{2}/v(F)^{2}+o(n^{2})$. The first such pair was found by Kovács and Nagy [14]: $F=C_{5}$ and $G=C_{3}$. Later, Balogh, Liebenau, Mattos and Morrison [1] refined this result by showing $(n^{2}-10n)/25\leq ex_{C_{5}}(n,C_{3})\leq(n^{2}+3n)/25+o(n)$. Characterizing when $ex_{F}(n,G)$ attains the lower bound is a challenging problem. Instead, we first study conditions that prevent it from attaining the lower bound.
@@ 4 Multicolor Turán number of cycles
    every edge in $E(C)$ lies between some consecutive parts $U_{i}$ and $U_{i+1}$, and
@@
$n\geq|W_{\ell}|\geq|W_{1}|\left(\frac{\delta}{16(k+1)^{2}\ell}\right)^{\ell-k}\geq\frac{\delta^{\ell-k+1}}{[16(k+1)^{2}\ell]^{\ell-k}},$
@@
To establish the foundation for the subsequent proof, we first define a rainbow common neighbor and introduce a classification of vertex pairs based on this notion. A vertex $w$ is called a rainbow common neighbor of two distinct vertices $x$ and $y$ if $w$ is adjacent to both $x$ and $y$, and the colors on the edges $wx$ and $wy$ are distinct. Based on the number of such neighbors, we classify an unordered pair of distinct vertices $\{u,v\}$ as follows: If $\{u,v\}$ has three rainbow common neighbors, it is called a large vertex pair; if $\{u,v\}$ has two rainbow common neighbors, it is called a medium vertex pair; if $\{u,v\}$ has one rainbow common neighbor, it is called a small vertex pair.
@@
Figure 2: Rainbow common neighbors of $u,v$. The second figure demonstrates the sharpness for the case $k=6$.
@@
Let $\{u,v\}$ be a large vertex pair and $x_{1},x_{2},x_{3}$ be their rainbow common neighbors. Then the coloring pattern is of Type 1. Furthermore, there exists additional two vertices $y_{1},y_{2}$ such that the union of black, green and red $C_{5}$s is depicted in Figure 3(1).

### 24 Mar 2026 v2

@@ Abstract
Motivated by the Ruzsa-Szemerédi problem, Imolay, Karl, Nagy, and Váli studied a variant of Turán number $ex_{F}(n,G)$ (called the $F$-multicolor Turán number of $G$), defined as the maximum number of edge-disjoint copies of $F$ on $n$-vertex set such that there is no copies of $G$ whose edges come from distinct copies of $F$. They proved that if there is no homomorphism from $G$ to $F$, then $n^{2}/v(F)^{2}+o(n^{2})\leq ex_{F}(n,G)\leq ex(n,G)/e(F)+o(n^{2})$, and otherwise $ex_{F}(n,G)=o(n^{2})$. The quantity $ex_{F}(n,G)$ asymptotically equals the maximum size of an $F$-packing in an $n$-vertex $G$-free graph, and attains the upper bound $ex(n,G)/e(F)+o(n^{2})$ if and only if $\chi(G)>\chi(F)$.
@@ 1 Introduction
The $(6,3)$-problem can be equivalently stated as follows: determine the maximum number of edge-disjoint triangles in an $n$-vertex graph such that no triangle is formed by taking one edge from each of three distinct triangles. Imolay, Karl, Nagy and Váli [13] generalized this problem by considering the following: given two fixed graphs $F$ and $G$, determine the maximum number of edge-disjoint copies of $F$ in an $n$-vertex graph such that there is no copy of $G$ whose edges come from distinct copies of $F$. This maximum, denoted by $ex_{F}(n,G)$, is called the $F$-multicolor Turán number of $G$. Clearly, if $F$ and $G$ are $r$-uniform hypergraphs, then $ex_{F}(n,G)=O(n^{r})$.
@@
In what follows, we restrict our discussion to simple graphs. Imolay, Karl, Nagy and Váli [13] proved that $ex_{F}(n,G)=\Theta(n^{2})$ if and only if there is no homomorphism from $G$ to $F$, and for the nondegenerate case ($ex_{F}(n,G)=\Theta(n^{2})$),
@@
Li [15] extended the above results to uniform hypergraphs. We say $ex_{F}(n,G)$ attains its lower bound (resp. upper bound) in (1) if $ex_{F}(n,G)=n^{2}/v(F)^{2}+o(n^{2})$ (resp. $ex_{F}(n,G)=ex(n,G)/e(F)+o(n^{2})$). Note that $\chi(G)>\chi(F)$ implies that there is no homomorphism from $G$ to $F$. Imolay, Karl, Nagy and Váli [13] showed that $ex_{F}(n,G)$ attains its upper bound when $\chi(G)>\chi(F)$. In fact, using the stability of extremal graphs for $ex(n,G)$, one can show that this condition is also necessary: $ex_{F}(n,G)$ attains its upper bound if and only if $\chi(G)>\chi(F)$ [15].
Imolay, Karl, Nagy and Váli [13] posed the question of characterizing all graph pairs $(F,G)$ for which $ex_{F}(n,G)=n^{2}/v(F)^{2}+o(n^{2})$. The first such pair was found by Kovács and Nagy [14]: $F=C_{5}$ and $G=C_{3}$. Later, Balogh, Liebenau, Mattos and Morrison [1] refined this result by showing $(n^{2}-10n)/25\leq ex_{C_{5}}(n,C_{3})\leq(n^{2}+3n)/25+o(n)$. Characterizing when $ex_{F}(n,G)$ attains the lower bound is a challenging problem. Instead, we first study conditions that prevent it from attaining the lower bound.
@@ 4 Multicolor Turán number of cycles
    every edge in $E(C)$ lies either in some $U_{i}$ or between consecutive parts $U_{i}$ and $U_{i+1}$, and
@@
$n\geq|W_{\ell}|\geq|W_{1}|\left(\frac{\delta}{8(k+1)^{2}\ell}\right)^{\ell-k}\geq\frac{\delta^{\ell-k+1}}{[8(k+1)^{2}\ell]^{\ell-k}},$
@@
To establish the foundation for the subsequent proof, we first define a rainbow common neighbor and introduce a classification of vertex pairs based on this notion. A vertex $w$ is called a rainbow common neighbor of two distinct vertices $x$ and $y$ if $w$ is adjacent to both $x$ and $y$, and the colors on the edges $wx$ and $wy$ are distinct. Based on the number of such neighbors, we classify an unordered pair of distinct vertices $\{u,v\}$ as follows: If $\{u,v\}$ has three rainbow common neighbors, it is called a large vertex pair; if $\{u,v\}$ has two rainbow common neighbors, it is called a medium vertex pair; if $\{u,v\}$ has at most rainbow common neighbor, it is called a small vertex pair.
@@
Figure 2: Rainbow common neighbors of $u,v$ (left side). The right side figure demonstrates the sharpness for the case $k=6$.
@@
Let $\{u,v\}$ be a large vertex pair and $x_{1},x_{2},x_{3}$ be their rainbow common neighbors. Then the coloring pattern is of Type 1. Furthermore, there exists additional two vertices $y_{1},y_{2}$ such that the union of black, green and red $C_{5}$s is depicted in Figure 3.